\begin{thebibliography}{999999}
\bibitem{ALL}
 C. Aistleitner, G. Larcher and M. Lewko. Additive energy and the Hausdorff dimension
of the exceptional set in metric pair correlation problems, with an appendix by J.
Bourgain, Israel J. Math., to appear.

\bibitem{BGS}
A.  Balog, A. Granville and J. Solymosi.
{\em Gaps between fractional parts, and additive combinatorics}. 
Quarterly Journal of Mathematics, 68 (2017), 1--12.


\bibitem{BBRR}
V. Blomer, J.  Bourgain, M. Radziwi\l\l \, and Z. Rudnick. Small gaps in the spectrum of the rectangular billiard.  	arXiv:1604.02413 [math.AP].  To appear in Annales Scientifiques de l'\'Ecole Normale Sup\'erieure.


\bibitem{BCGW}
T. F. Bloom, S. Chow, A. Gafni and  A. Walker. Additive energy and the metric Poissonian property. 	arXiv:1709.02634 [math.NT]

\bibitem{BS}
A. Bondarenko and K. Seip. GCD sums and complete sets of square-free numbers.
Bull. Lond. Math. Soc., 47(1):29--41, 2015.

%\bibitem[C]{C}
%D. Carmon, Evenly divisible rational approximations of quadratic irrationalities. 	arXiv:1612.00382 [math.NT]
%to appear in the Israel Jour. of Math. 

\bibitem{DH}
T. Dyer and G. Harman. Sums involving common divisors, J. London Math. Soc. 34 (1986) 1--11. 

\bibitem{HB}
D.R. Heath-Brown. Pair correlation for fractional parts of $\alpha n^2$. Math. Proc. Cambridge Philos. Soc. 148 (2010), no. 3, 385--407.

\bibitem{LT}
T. Lachmann and N. Technau, On Exceptional Sets in the Metric Poissonian Pair
Correlations Problem, arXiv:1708.08599.

\bibitem{Levy}
P.  L\'evy. {\em Sur la division d'un segment par des points choisis au
hasard}, C.R. Acad. Sci. Paris 208 (1939), 147--149.

%\bibitem{LR}
%M. Lewko and M. Radziwi\l\l. 
%Refinements of G\'al's theorem and applications. Adv. Math. 305 (2017), 280--297. 


\bibitem{MS}
J. Marklof and A. Str\"ombergsson. Equidistribution of Kronecker sequences along closed horocycles. Geom. Funct. Anal. 13 (2003), 1239--1280.

%\bibitem{3gap}
% G. Polanco, D. Schultz, Daniel and A. Zaharescu. 
%Continuous distributions arising from the three gap theorem. Int. J. Number Theory 12 (2016), no. 7, 1743--1764. 


\bibitem{RS}
 Z. Rudnick and P. Sarnak. The pair correlation function of fractional parts of
polynomials. Comm. Math. Phys., 194(1):61--70, 1998.


\bibitem{RZ}
 Z. Rudnick and A. Zaharescu. A metric result on the pair correlation of fractional
parts of sequences. Acta Arith., 89(3):283--293, 1999.

\bibitem{VZ}
M. V\^aj\^aitu and A. Zaharescu. {\em Distinct gaps between fractional parts of sequences}. 
Proc. Amer. Math. Soc. 130 (2002), 3447--3452.

\bibitem{Walker} A. Walker. The primes are not metric Poissonian, arXiv:1702.07365.

\end{thebibliography}

\documentclass[10pt]{amsart}
\usepackage{mathrsfs,amssymb}
\usepackage{graphicx}
\begin{document}

\newtheorem{theorem}{Theorem}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{conjecture}[theorem]{Conjecture}
\newtheorem{question}[theorem]{Question}
\newtheorem{problem}[theorem]{Problem}
\theoremstyle{definition}
\newtheorem{definition}{Definition}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}

\renewcommand{\labelenumi}{(\roman{enumi})}
\def\theenumi{\roman{enumi}}


\renewcommand{\Re}{\operatorname{Re}}
\renewcommand{\Im}{\operatorname{Im}}

\def\scrA{{\mathcal A}}
\def\scrB{{\mathcal B}}
\def\scrD{{\mathcal D}}
\def\scrL{{\mathcal L}}
\def\scrS{{\mathcal S}}

\def \G {{\Gamma}}
\def \g {{\gamma}}
\def \R {{\mathbb R}}
\def \C {{\mathbb C}}
\def \Z {{\mathbb Z}}
\def \Q {{\mathbb Q}}
\def \TT {{\mathbb T}}
\newcommand{\T}{\mathbb T}
\def \GinfmodG {{\Gamma_{\!\!\infty}\!\!\setminus\!\Gamma}}
\def \GmodH {{\Gamma\setminus\H}}
\def \vol {\hbox{vol}}
\def \sl  {\hbox{SL}_2(\mathbb Z)}
\def \slr  {\hbox{SL}_2(\mathbb R)}
\def \psl  {\hbox{PSL}_2(\mathbb R)}

\newcommand{\mattwo}[4]
{\left(\begin{array}{cc}
                        #1  & #2   \\
                        #3 &  #4
                          \end{array}\right) }

\newcommand{\rum}[1] {\textup{L}^2\left( #1\right)}
\newcommand{\norm}[1]{\left\lVert #1 \right\rVert}
\newcommand{\abs}[1]{\left\lvert #1 \right\rvert}
\newcommand{\inprod}[2]{\left \langle #1,#2 \right\rangle}

\renewcommand{\i}{{\mathrm{i}}}

\newcommand{\area}{\operatorname{area}}
\newcommand{\ecc}{\operatorname{ecc}}

\newcommand{\Op}{\operatorname{Op}}
\newcommand{\dom}{\operatorname{Dom}}
\newcommand{\Dom}{\operatorname{Dom}}
\newcommand{\tr}{\operatorname{tr}}

\newcommand{\Norm}{\mathcal N}
\newcommand{\simgeq}{\gtrsim}%
\newcommand{\simleq}{\lesssim}

\newcommand{\UN}{U_N}
\newcommand{\OPN}{\operatorname{Op}_N}
\newcommand{\HN}{\mathcal H_N}
\newcommand{\TN}{T_N}  
\newcommand{\PDO}{\Psi\mbox{DO}}

\newcommand{\intinf}{\int_{-\infty}^\infty}
\newcommand{\lcm}{\operatorname{lcm}}
\newcommand{\E}{\mathbb E}
\newcommand{\Prob}{\operatorname{Prob}}
\newcommand{\Var}{\operatorname{Var}}
\newcommand{\conv}{*}

\title{A metric theory of minimal gaps}
\author{Ze\'ev Rudnick}
\address{Raymond and Beverly Sackler School of Mathematical Sciences,
Tel Aviv University, Tel Aviv 69978, Israel}
\email{rudnick@post.tau.ac.il}

\date{\today}

\begin{abstract}
We study the minimal gap statistic for fractional parts of sequences of the form $\mathcal A^\alpha = \{\alpha a(n)\}$ where $\mathcal A = \{a(n)\}$ is a sequence of distinct of integers. Assuming that the additive energy of the sequence is close to its minimal possible value, we show that for almost all $\alpha$, the minimal gap $\delta_{\min}^\alpha(N)=\min\{\alpha a(m)-\alpha a(n)\bmod 1: 1\leq m\neq n\leq N\}$ is close to that of a random sequence. \end{abstract}

\maketitle

We start with a sequence of points $\mathcal X=\{x_n:n=1,2,\dots\}\subset \R/\Z$ in the unit interval/circle, which we assume is asymptotically uniformly distributed: 
For any subinterval $\subset \R/\Z$, we have 
\begin{equation}\label{UD}
\lim_{N\to \infty} \frac 1N\#\{n\leq N: x_n\in I\} = |I| \;.
\end{equation}
In particular, the mean spacing between the points lying in any subinterval is $1/N$. Our goal is to understand the {\em minimal gap} 
$$ \delta_{\min}(\mathcal X,N) = \min(|x_n-x_m|: n,m\leq N, n\neq m)$$
(with  a suitable  modification for wrapping around). 

For {\em random} points,  namely  $N$ independent uniform points in the unit interval (Poisson process), the minimal gap is almost surely of size $1/N^2$ \cite{Levy}. In this note we study the metric theory of the minimal gap statistic for a class of deterministic sequences of fractional parts, such as fractional parts of polynomials. The case of quadratic polynomials $x_n =\alpha n^2$ has its roots in the recent paper \cite{BBRR}, which studies the more complicated case of the minimal gap statistic for the sequence of eigenvalues of the Laplacian on a rectangular billiard, namely the points 
$\{\alpha m^2+n^2 : m,n\geq 1\}$ on the real line. 

We fix   a sequence   $\mathcal A=\{a(n):n=1,2,\dots\}\subset \mathbb Z$ of distinct integers ($a(n)\neq a(m)$ if $m\neq m$), and study the minimal gap statistic of fractional parts of the set 
$$\mathcal A^\alpha = \{\alpha a(n) \bmod 1:n=1,2, \dots\}\subset \R/\Z \;.
$$
(it is an old result of Weyl that $\mathcal A^\alpha$ satisfies \eqref{UD} for almost all $\alpha$). 
We want to know under which conditions we can show that for {\em almost all} $\alpha$, the minimal gap statistics 
$$\delta_{\min}^\alpha(N)=\delta_{\min}( \mathcal A^\alpha,N)$$ 
follows that of the random case, that is of size about $1/N^2$ for almost all $\alpha$.  
It is easy to see that we cannot have much smaller minimal gaps:
\begin{theorem}\label{thm delta big}
Assume that $\mathcal A$ consists of distinct integers.  Then for all  $\eta>0$, for almost all $\alpha$, 
$$
\delta_{\min}^\alpha(N)>\frac 1{N^{2+\eta}},\quad \forall N>N_0(\alpha) \;.
$$
\end{theorem}

To make the minimal gap small, we give a criterion in terms of the ``additive energy" $E(\mathcal A,N) $ of the sequence: 
$$
E(\mathcal A,N) :=\#\{(n_1,n_2,n_3,n_4)\in [1,N]^4: a(n_1)+a(n_2) = a(n_3)+a(n_4) \}  \;.
$$ 
Note that $N^2\leq E(\mathcal A,N) \leq N^3$. 
The result is
\begin{theorem}\label{thm min}
Assume that $\mathcal A$ consists of distinct integers, and that the additive energy satisfies
$$ E(\mathcal A,N)\ll N^{2+o(1)},\quad \forall N\gg 1\;.
$$
Then  for all $\eta>0$, for almost all $\alpha$, 
$$
      \delta_{\min}^\alpha(N)<\frac 1{N^{2-\eta}},\quad \forall N>N_0(\alpha) \;.
$$
\end{theorem}
 Examples:  
For $a(n)=n^d$, $d\geq 2$, it is shown in \cite{RS} that the $E(\mathcal A,N)\ll N^{2+o(1)}$. For lacunary sequences, we have   $E(\mathcal A,N)\ll N^2$  \cite{RZ}.  Hence Theorem~\ref{thm min} applies to these sequences.

Relaxing the required bound on the additive energy will give a weaker result on the minimal spacing, basically that $\delta_{\min}^{\alpha}(N)<E(\mathcal A,N)/N^{4-\eta}$ almost surely. For this to be nontrivial, we need the additive energy to be no bigger than $E(\mathcal A,N)\ll N^{3-\eta'}$ for some $\eta'>0$. A notable case where the additive energy is bigger is that of $\mathcal A=\mathcal P$ being the sequence of primes, where $E(\mathcal P,N) \approx N^3/\log N$.  
In this case, we cannot have gaps much larger than the average gap: A simple argument shows that given any $\varepsilon>0$, for almost all $\alpha$, we have $\delta_{\min}(\mathcal P^\alpha,N)\gg 1/(N(\log N)^{2+\varepsilon})$, see \S~\ref{sec:dense}.



\noindent{\bf Acknowledgements:} We thank Andrew Granville and Niclas Technau for their comments. 
The work was supported  by the European Research Council under the European Union's Seventh
Framework Programme (FP7/2007-2013)/ERC grant agreement n$^{\text{o}}$ 320755.   
 
\section{A bilinear statistic }

To study the minimal gap, we introduce a statistics counting all possible gaps: 
We start with a smooth, compactly supported window function $f\in C_c^\infty([-1/2,1/2])$, which is non-negative: $f\geq 0$, 
and of unit mass $\int f(x)dx=1$, and define 
$$F_M(x) = \sum_{j\in \Z} f(M(x+j))$$
which is localized on the scale of $1/M$, and periodic: $F_M(x+1) = F_M(x)$. We then set 
$$
D_{\mathcal A}(N,M)(\alpha) = \sum_{\substack{1\leq m,n\leq N\\m\neq n}} F_M(\alpha a(n)-\alpha a(m)) \;.
$$

The expected value of $D_{\mathcal A}(N,M)$ is easily seen to equal 
\begin{equation}\label{E(D)} 
\int_0^1 D_{\mathcal A}(N,M)(\alpha)d\alpha = \frac{N(N-1)}{M}\sim \frac{N^2}{M}\;.
\end{equation}
This already suffices to show that minimal gaps cannot typically be small (Theorem~\ref{thm delta big}), see \S~\ref{sec:lower}.

We will bound the  variance of $D_{\mathcal A}(N,M)$, from which Theorem~\ref{thm min} will follow: 
\begin{proposition}\label{prop var}
\begin{equation*}
\Var D_{\mathcal A}(N,M) \ll \frac{1}{M} N^{\epsilon}E(\mathcal A,N)  \;.
\end{equation*}
\end{proposition}

The statistic $D_{\mathcal A}(N,M)(\alpha)$ is related to the pair correlation function of the sequence $\mathcal A^\alpha$, which in our notation is $D_{\mathcal A}(N,N)(\alpha)/N$. Pair correlation measures gaps on the scale of the mean spacing, assumed here to be $1/N$, corresponding to $M=N$, here we are looking at much smaller scales of $M$ close to $N^2$. 

The metric theory of the pair correlation function of fractional parts was initiated in \cite{RS}, where the sequences $a(n) = \alpha n^d$ were shown to almost surely have Poissonian pair correlation for $d\geq 2$ (see \cite{MS, HB} for different proofs of the quadratic case $d=2$). The problem has since been studied in several other cases  and  has recently been revived   in an abstract setting \cite{ALL, Walker,  LT, BCGW}. In particular, a convenient criterion for almost sure Poisson pair correlation has been formulated by Aistleitner,  Larcher and Lewko \cite{ALL} in terms of the additive energy $E(\mathcal A,N)$ of the sequence. 
The proof of Proposition~\ref{prop var} is close to that of the analogous statement for the pair correlation function in \cite{ALL}, which in turn is based on \cite{RS, RZ}. 

\begin{proof}


The Fourier expansion of $F_M(x)$ is 
\begin{equation}
F_M(x) = \sum_{k\in \Z} \frac 1M \widehat{f}(\frac kM) e(kx)
\end{equation}
where $\widehat{f}(y) = \int_{-\infty}^\infty f(x)e^{-2\pi i xy }dx$. Inserting into the definition of $D(N,M)$ gives 
\begin{equation}
D(N,M)(\alpha) =\sum_{k\in \Z}  \frac 1M \widehat{f}(\frac kM)  \sum_{\substack{1\leq m,n\leq N\\m\neq n}}  e(k\alpha(a(m)-a(n)) \;.
\end{equation}
Integrating over $\alpha$ gives the expected value (we assume $a(m)\neq a(n)$ if $n\neq m$) 
$$
\int_0^1 D(N,M)(\alpha)d\alpha =\widehat{f}(0)\frac 1M N(N-1) =\frac{N(N-1)}{M} \;.
$$

The variance is the second moment of the sum over nonzero frequencies:
$$
\Var D(N,M) = \int_0^1 \Big| \sum_{0\neq k\in Z}  \frac 1M \widehat{f}(\frac kM)  \sum_{\substack{1\leq m,n\leq N\\m\neq n}}  e(k\alpha(a(m)-a(n)) \Big|^2 d\alpha \;.
$$
Squaring out and integrating gives 
\begin{multline*}
\Var D(N,M) = 
 \sum_{k_1,k_2\neq 0} \frac 1{M^2}  \widehat{f}(\frac {k_1}M) \widehat{f}(\frac {k_2}M) \times 
\\
 \#\{m_1\neq n_1, m_2\neq n_2: k_1(a(m_1)-a(n_1)) = k_2 a(m_2)-a(n_2))\} \;.
\end{multline*}
We now follow \cite[Lemma 3]{ALL} to convert this to ``GCD sums". Let 
$$
R(v) = \#\{1\leq m\neq n\leq N: a(m)-a(n) = v\} \;.
$$ 
Then 
$$
\Var D(N,M) = \sum_{ v_1,v_2\neq 0} \sum_{k_1,k_2\neq 0} c(k_1)c(k_2)  R(v_1)R(v_2)\delta(k_1v_1=k_2v_2)
$$
where we set 
$$ c(k) :=\frac 1M \widehat{f}(\frac kM) \;.
$$

\begin{lemma}
Let $f\in C_c^\infty(\R)$. For any nonzero integers $v_1, v_2\neq 0$, 
\begin{equation}\label{improved gcd}
\sum_{k_1,k_2\neq 0} c(k_1)c(k_2)   \delta(k_1v_1=k_2v_2)
\ll_f   \frac 1M \frac{\gcd(v_1,v_2)}{\sqrt{|v_1v_2|}} 
\;.
\end{equation}
\end{lemma}
\begin{proof}
For $k_1,k_2\neq 0$, 
we have $k_1v_1=k_2 v_2$ if and only if 
$$(k_1,k_2) = \ell (\frac{v_2}{\gcd(v_1,v_2)}, \frac{v_1}{\gcd(v_1,v_2)})
$$
for some nonzero integer $0\neq \ell \in \Z$. Abbreviating $a_i  = v_i/(M\gcd(v_1,v_2))$, we find  
\begin{equation*}
\begin{split}
\sum_{k_1,k_2\neq 0} c(k_1)c(k_2)\delta(k_1v_1=k_2v_2) 
&= \sum_{0\neq \ell\in \Z} c(\ell \frac{v_2}{\gcd(v_1,v_2)} ) c(\ell \frac{v_1}{\gcd(v_1,v_2)}) 
 \\& 
= \frac 1{M^2}  \sum_{0\neq \ell\in \Z}\widehat{f}(a_2\ell ) \widehat{f}(a_1 \ell ) 
\end{split}
\end{equation*}
so that it suffices to show that 
\begin{equation}\label{improved gcd2}
  \sum_{0\neq \ell\in \Z}\widehat{f}(a_1 \ell  ) \widehat{f}(a_2 \ell  ) \ll_f \frac 1{\sqrt{|a_1a_2|}} = \frac{M\gcd(v_1,v_2)}{\sqrt{|v_1v_2|}} \;.
\end{equation}

Applying Cauchy-Schwarz   we get 
$$
\sum_{0\neq \ell\in \Z}\widehat{f}(\ell a_1) \widehat{f}(\ell a_2 ) \leq \Big(\sum_{\ell\neq 0} \widehat{f}(a_1\ell)^2\Big)^{1/2} \Big(\sum_{\ell \neq 0} \widehat{f}(a_2\ell)^2\Big)^{1/2} \;.
$$
We will obtain \eqref{improved gcd2} if we show that for any $a>0$, 
$$
\sum_{\ell \neq 0} \widehat{f}(a\ell)^2 \ll_f \frac 1 a  \;.
$$
Indeed, if $0<a\ll 1$ then we get a Riemann sum for $(\widehat{f}\;)^2$:
$$
\sum_{\ell \neq 0} \widehat{f}(a\ell)^2\sim \frac 1a \intinf \widehat{f}(y)^2 dy =  \frac 1a \intinf f(x)^2 dx \;.
$$
If $ a\gg 1$ then use the decay rate of the Fourier transform: For $y\neq 0$, 
$$|\widehat{f}(y)|\leq \frac 1{2\pi |y|}\intinf|f'(x)|dx \;,
$$
to obtain
$$
\sum_{\ell \neq 0} \widehat{f}(a\ell)^2  \ll  \sum_{\ell\neq 0} (\frac {\intinf |f'(x)|dx}{ |a\ell|})^2 \ll_f \frac 1{a^2}
$$
which for $a\gg 1$ is $\ll 1/a$. 
\end{proof}

   Hence 
$$
\Var D(N,M)\ll     \frac 1M\sum_{ v_1,v_2\neq 0}  R(v_1)R(v_2)\frac{\gcd(v_1,v_2)}{\sqrt{|v_1v_2|}} \;.
$$
According to the GCD bounds of \cite{DH},   
$$
\sum_{ v_1,v_2\neq 0}  R(v_1)R(v_2)\frac{\gcd(v_1,v_2)}{\sqrt{|v_1v_2|}}  \ll   \exp(\frac{10 \log N}{\log\log N})
\sum_v R(v)^2  
$$
(see \cite{BS} for an essentially optimal refinement).
Now 
$$ \sum R(v)^2 = \#\{m_i,n_j\leq N, m_1\neq n_1, m_2\neq m_2 : a(m_1)-a(n_1)=a(m_2)-a(n_2)\}
$$
is at most  the additive energy $E(\mathcal A,N)$. Thus 
$$ \Var D(N,M)\ll \frac{1}{M} N^{\epsilon}E(\mathcal A,N)$$
as claimed. 
\end{proof}

\begin{corollary}\label{almost sure D}
Assume that the additive energy satisfies $E(\mathcal A,N)<N^{2+o(1)}$. If $M<N^{2-\eta}$,  then for almost all $\alpha$, 
$$ D_{\mathcal A}(N,M)(\alpha)\sim \frac{N^2}{M} \;.$$
\end{corollary}

\begin{proof}
Take $N_k = \lfloor k^{4/\eta }\rfloor$, so that $\sum_k N_k^{-\eta/2}<\infty$, 
and pick any $M_k <N_k^{2-\eta}$, then  we find from Proposition~\ref{prop var}
\begin{equation*}
\begin{split}
 \sum_k \int_0^1 \Big| \frac{D(N_k,M_k)(\alpha)}{N_k(N_k-1)/M_k}-1 \Big|^2 d\alpha &= 
\sum_k \frac{\Var D(N_k,M_k)}{ (N_k(N_k-1)/M_k)^2} 
\\
&<\sum_k N_k^{o(1)} \frac{E(N_k)M_k}{N_k^4}\ll
\sum_k \frac 1{N_k^{ \eta/2} }<\infty
\end{split}
\end{equation*}
and so for almost all $\alpha$
\begin{equation}\label{ae for D_k}
D(N_k,M_k)(\alpha) \sim \frac{N_k^2}{M_k}, \quad \forall k>k_0(\alpha)  \;.
\end{equation}
Apriori the set depends on the test function $f$, but that can be taken care of by a standard diagonalization procedure;  for our purposes we only need one test function.

Given $N\gg 1$, there is a unique value of $k$ so that $N_k\leq N<N_{k+1}$. 
Note that $N/N_k = 1+O(N^{-\eta/4})$. 
Since $M<N^{2-\eta}\sim N_k^{2-\eta}<N_{k+1}^{2-\eta}$, we know   \eqref{ae for D_k} that almost surely $D(N_k,M)/(N_k^2/M) \to 1$. 

Note that 
\begin{equation*}\label{lem:monotonicity}
D(N_k, M)\leq D(N,M)  \leq D(N_{k+1},M) \;.
\end{equation*}
This is because the sums $D(N,M)$ consist of non-negative terms, and hence 
\begin{multline*}
D(N,M) = \sum_{1\leq m\neq n\leq N} F_M(\alpha(a(m)-a(n))) 
\\
\geq  \sum_{1\leq m\neq n\leq N_k} F_M(\alpha(a(m)-a(n))) =D(N_k,M)
\end{multline*}
(we dropped all pairs $(m,n) $ where $\max(m,n)>N_k$).

Since  $N/N_k = 1+O(N^{-\eta/4})$, we have  
$$ 
\frac{D(N_k,M)}{N_k^2/M} \leq \frac{D(N,M) }{N^2/M(1+O(N^{-\eta/4}))} = \frac{D(N,M)}{N^2/M}(1+O(N^{-\eta/4}) )
$$
and likewise 
$$ \frac{D(N,M)}{N^2/M}\leq \frac{D(N_{k+1},M)}{N_{k+1}^2/M} (1+O(N^{-\eta/4})) \;.
$$
Since we know that almost surely $D(N_k,M)/(N_k^2/M) \to 1$, we deduce that almost surely also $ D(N,M)/(N^2/M)\to 1$. 
\end{proof}

\begin{corollary}
 Theorem~\ref{thm min} holds.
\end{corollary}
\begin{proof}
Fix $\eta>0$, and let $M=\frac 12 N^{2-\eta}$. Since $D(N,M)(\alpha) \sim \frac{N^2}{M} >1$ by Corollary~\ref{almost sure D}, 
we have a gap of size at most $1/(2M)=1/N^{2-\eta}$, that is $\delta_{\min}^\alpha(N)<1/N^{2-\eta}$ almost surely. 
\end{proof}

\section{Lower bounds: Proof of Theorem~\ref{thm delta big}}\label{sec:lower}
We take any sequence of integers $\mathcal A = \{a(n)\}$ with distinct elements. We want to show that for any $\eta>0$, almost surely, 
$$\delta^\alpha(N)>1/N^{2+\eta}, \quad \forall N>N_0(\alpha)\;.
$$
Let $N_k = \lfloor k\rfloor^{2/\eta}$. We claim that it suffices to show that for almost all $\alpha$, 
\begin{equation}\label{delta small}
\delta_{\min}^\alpha(N_k)>2/N_k^{2+\eta}, \quad \forall k>k_0(\alpha)\;.
\end{equation}

Indeed, note that if $N_k\leq N<N_{k+1}$ then $\delta_{\min}^\alpha(N)\geq \delta_{\min}^\alpha(N_{k+1})$. 
Since $N_{k+1}\sim N$, by \eqref{delta small} we have for almost all $\alpha$  
$$ \delta_{\min}^\alpha(N)\geq \delta_{\min}^\alpha(N_{k+1})>2/N_{k+1}^{2+\eta}>1/N^{2+\eta}$$
for $N>N_0(\alpha)$.

To prove \eqref{delta small}, it suffices,  by the Borel-Cantelli lemma,  to show that 
\begin{equation}\label{sum prob delta}
\sum_k \Prob\Big(\delta_{\min}^\alpha(N_k)\leq 2/N_k^{2+\eta} \Big) <\infty\;.
\end{equation}

In the definition of $D(N,M)=D_f(N,M)$, choose $f$ so that $f(x)\geq 1$ if $|x|\leq 1/4$ (and in addition, $f\geq 0$ is non-negative, $\intinf f(x)dx=1$, $f$ is smooth and supported in $[-1/2,1/2]$). 
Now note that for such $f$, if $D_f(N,M)<1$ then $\delta_{\min}^\alpha(N)>1/(4M)$. This is because $D_f(N,M)$ is a sum of non-negative terms, and if there is one gap of size $\leq 1/(4M)$ then the corresponding term $F_M(\alpha(a(m)-a(n)) = \sum_j f(M(\alpha(a(m)-a(n))+j)\geq 1$ by the choice of $f$, so that $D_f(N,M)\geq 1$. Thus we find that 
$$
\delta_{\min}^\alpha(N)\leq \frac 1{4M} \quad  \Rightarrow \quad D_f(N,M)\geq 1 
$$
and hence 
\begin{equation}\label{bd for prob delta}
\Prob\Big(\delta_{\min}^\alpha(N )\leq \frac 1{4M}\Big) \leq \Prob\Big( D_f(N, M ) \geq 1\Big) \;.
\end{equation}

Now since $D_f\geq 0$, 
$$
\Prob\Big( D_f(N ,M ) \geq  1 \Big) \leq \int_0^1 D_f(N , M)(\alpha)d\alpha
$$
so that by \eqref{E(D)}  for $M_k=\dfrac 18 N_k^{2+\eta}$, 
$$  \int_0^1 D(N_k,M_k)(\alpha)d\alpha \sim 8N_k^{-\eta}\ll \frac 1{k^2}$$
which together with \eqref{bd for prob delta} proves \eqref{sum prob delta}, hence \eqref{delta small}. 
This proves Theorem~\ref{thm delta big}.

\section{Minimal gaps for the primes}\label{sec:dense}

Let $a(n) = p_n$, the $n$-th prime. By Khinchin's theorem, for all $\varepsilon>0$  there is a set of full measure of $\alpha$'s so that 
$|| q\alpha ||\geq 1/(q(\log q)^{1+\varepsilon})$ for any integer $q\geq q_0(\alpha)$.  
In particular, for such $\alpha$, the gap between fractional parts of $\alpha p_n \bmod 1$ are
$$ ||\alpha(p_m-p_n)|| \gg \frac 1{|p_m-p_n|(\log |p_m-p_n|)^{1+\varepsilon}} \geq \frac 1{ N(\log N)^{2+\varepsilon}} \;,
 $$
since $|p_m-p_n|\leq p_N\sim N\log N$ for $m<n\leq N$. Hence  for such $\alpha$, the minimal gap satisfies 
$\delta_{\min}^\alpha(N)>1/N( \log N)^{2+\varepsilon}$. 

A similar argument applies to other dense cases, such as the sequence of squarefree integers.  
An extreme case is that when $\mathcal A=\mathbb N$ is the sequence of all natural numbers. The argument above gives the minimal gap here is, for almost all $\alpha$, at least $\delta_{\min}^\alpha(N)\gg 1/(N(\log N)^{1+\varepsilon})$. 
Note that in this case the ``three-gap" theorem shows that there are at most three distinct gaps between the fractional parts 
$\{\alpha n\bmod 1:n\leq N\}$. 
Concerning other ``dense" sequences, it is known that for any sequence of integers $\mathcal A \subset [1,M]$, the  fractional parts 
$\alpha a(m)\bmod 1$ have at most $O(\sqrt{M})$ distinct gaps \cite{VZ, BGS}.

\end{document}